You do 1000 J of work lifting a 25 kg box from the floor to a shelf, only to have it accidentally fall off. About how high is the shelf?
Answer #1

Work-done is given by:

W = F.d = F*d*cos A

F = Force applied by you against gravity = m*g

m = mass of box = 25 kg

d = displacement of box = distance from floor to shelf = ?

A = Angle between Force and displacement = 180 deg

W = Work-done by you against gravity = -1000 J (-ve sign means work-done against gravity)

So,

d = W/(F*cos A)

d = -1000/(25*9.8*cos 180 deg)

d = height of shelf = 4.08 m

Please Upvote.